Use technology to find the​ P-value for the hypothesis test described below.

The claim is that for the population of adult​ males, the mean platelet count is

μ<211 . The sample size is n=52 and the test statistic is t=−1.738.

​P-value=____________

​(Round to 3 decimall places
Answer #1

Solution:

This is the left tailed test

n = 52

df = n - 1 = 52 - 1 = 51

P-value = P(t < -1.738) = 0.044